OpenStudy (aglna):

The volume of a cone is 62.8 cubic inches the height of the cone is 15 inches wha it's the radius of the cone my answer 8 inches 2 inches 4inches 8inches 9inches

OpenStudy (mathstudent55):

\(\Large V_{cone} = \dfrac{1}{3} \pi r^2h\)

OpenStudy (mathstudent55):

Plug in the given values and solve for r.

OpenStudy (mathstudent55):

I don't get 8 inches.

OpenStudy (wolf1728):

Cone Volume = (π • r² • h) ÷ 3 3 * Volume / (PI * h) = r² 3 * 62.8 / (3.14159 * 15) 188.4 / 47.124 = r² 4 = r² So r = ??

OpenStudy (aglna):

2 inches

OpenStudy (wolf1728):

That's right Aglna !!

OpenStudy (mathstudent55):

Correct.

OpenStudy (aglna):

Thanks for the help guys!!

OpenStudy (wolf1728):

u r welcome

OpenStudy (mathstudent55):

You're better off learning how to solve the problem than guessing. With a guess you have just a 25% probability of being correct.